OpenStudy (anonymous):

a number plus the reciprocal of the number is 5/4 of the number. What is the number? (fyi answer is -2 or 2)

OpenStudy (watchmath):

\(x+\frac{1}{x}=\frac{5}{4}x\) \(\frac{1}{x}=\frac{x}{4}\) \(4=x^2\) \(x=\pm 2\)

OpenStudy (anonymous):

how did u go from the 1st line to the 2nd line?

OpenStudy (watchmath):

subtract x from both sides